OpenStudy (anonymous):

Put the equation below in the form ax2 + bx + c = 0. x^2-5=-4x^2+3x help please!

OpenStudy (anonymous):

x^2 - 5=-4x^2 + 3x +4x^2 +4x^2 5x^2 - 5 = 3x -3x -3x 5x^2 - 3x - 5 = 0

OpenStudy (anonymous):

First you add 4x^2 to both sides. Then you subtract 3x from both sides.

OpenStudy (anonymous):

thank you sir :D